Exercise 6-4 Basic Segmented Income Statement [LO6-4] Royal Lawncare Company produces and sells two packaged products—Weedban and Greengrow. Revenue and cost information relating to the products follow: Product Weedban Greengrow Selling price per unit $ 8.00 $ 38.00 Variable expenses per unit $ 3.00 $ 10.00 Traceable fixed expenses per year $ 134,000 $ 34,000 Common fixed expenses in the company total $106,000 annually. Last year the company produced and sold 41,500 units of Weedban and 25,000 units of Greengrow. Required: Prepare a contribution format income statement segmented by product lines.
Answer #1

Answer

  • Contribution margin Income Statement

Total Company

Weedban

Greengrow

Sales

$1,282,000

$332,000

$950,000

Variable expenses

$539,500

$124,500

$415,000

Contribution margin

$742,500

$207,500

$535,000

Traceable Fixed expenses

$168,000

$134,000

$34,000

Product Line Segment Margin

$574,500

$73,500

$501,000

Common Fixed expenses not traceable to products

$106,000

Net Operating Income

$468,500

--Working

Weedban

Greengrow

Sales

=41500*8

=25000*38

Variable expenses

=41500*3

=41500*10

Contribution margin

$207,500

$535,000

Traceable Fixed expenses

134000

34000

Product Line Segment Margin

$73,500

$501,000